OpenStudy (anonymous):

Need help! Solve the system using substitution method c(n)=-3n-6 c(n)=n-4

OpenStudy (king):

n=-2

OpenStudy (king):

?

OpenStudy (king):

oh no sry!

OpenStudy (lalaly):

from the second equation n=c+4 substitute c+4 in n in the first equation c=-3(c+4)-6 solve for c then what u get for c u substitute it in any of the equations to find n

OpenStudy (king):

hw is n=c+4?

OpenStudy (king):

is that n*c?

OpenStudy (lalaly):

no its not, its c(n) like f(x) i guess

OpenStudy (king):

i know that wht i thought...but hw is n=c+4?

OpenStudy (lalaly):

c is the same as c(n)

OpenStudy (mimi_x3):

When you move the negative to the other side, its positive.

OpenStudy (king):

oh ok i didnt know that c is the same as c(n)

OpenStudy (lalaly):

or u can just subsitute c=n-4 and find n then find c

OpenStudy (king):

oh so i got c=-4.5

OpenStudy (king):

so n=8.5

OpenStudy (anonymous):

n=-0.5

OpenStudy (anonymous):

c(n)=-3n-6 c(n)=n-4 is this a elementary math? suppose c(n) = y and n = x, then the equation will be : y=-3x-6 ...(1) y=x-4...(2) arranging x from the equation (1) to the left side : y+3x=-6 ...(1) y=x-4 ... (2) subtitute y from (2) to (1) : (x-4) +3x = -6 solve this : 4x-4=-6 4x=-2 x= - .5 to get the value of y, subtitue this x to equation (2) : y=x-4 y=(-.5)-4 y=-4.5 that's it! the solution is y=-4.5 and x==.5 or c(n)=-4,5 and n=-.5

OpenStudy (anonymous):

that's Equations Not system function.....

OpenStudy (anonymous):

囧~